(Part 1)You are about to invest $15,000 into a project that will generate $5,500 of cash...

  1. (Part 1)You are about to invest $15,000 into a project that will generate $5,500 of cash flows each year for the next 3 years. If your cost of capital is 11%, then the present value of future cash flows is:
    1. -$1,559
    2. $1,559
    3. $1,029
    4. -$1,029
  1. (Part 2)The present value of the following cash flow stream at an interest rate of 10% per year is closest to:

$0 at Time 0;

$1,500 at the end of Year 1;

$3,000 at the end of Year 2;

$4,500 at the end of Year 3

  1. $6,934
  2. $3,559
  3. $2,329
  4. $7,329
Answer #1

Present value of this cash flow = -15000 + 5500 * (1 – 1.11^-3)/0.11 = -1559. Hence select option A

Present value of this cash flow = 1500(P/A, 10%, 3) + 1500(P/G, 10%, 3) = 1500*2.4869 + 1500*2.3291 = 7223. Hence select option D.